### v1

# Qubit-assisted quantum metrology

Abstract
We propose a quantum metrology protocol based on a two-step joint evolution of the probe system and an ancillary qubit and a single-shot projective measurement. With an optimized initialization of the ancillary qubit, the quantum Fisher information (QFI) about the phase parameter encoded in the probe system is found to be determined by the expectation value of the square of a time-optimized phase generator, independent of the probe state. Therefore, QFI can approach the Heisenberg scaling $N^{2}$ with respect to the quantum number $N$, even when the probe system is prepared in a classical state. We find that this scaling behavior is robust against the imperfections in preparing the ancillary qubit and controlling the evolution time. Using the time-reversal strategy, the classical Fisher information (CFI) in our metrology protocol is saturated with its quantum counterpart. Our work thus paves an economical way to realize the Heisenberg-scaling limit in metrology precision with no use of entanglement or squeezing.
I Introduction
Quantum metrology aims to use various quantum resources including entanglement and squeezing to excess the standard quantum limit (SQL), where the uncertainty decreases with $1/\sqrt{N}$, on performing measurements of physical parameters Sun et al. (2010); Ma et al. (2011); Genoni et al. (2012); Escher et al. (2012); Zhong et al. (2013). Examples of these parameters are the phases of evolving quantum systems encoded in multiple applications Paris (2009); Helstrom (1969); Yurke et al. (1986), such as gravitational wave detection Caves (1981), biological sensing Taylor and Bowen (2016); Mauranyapin et al. (2017), atomic clocks Ludlow et al. (2015); Katori (2011), and magnetometry Jones et al. (2009). A conventional protocol for precision metrology reaching the Heisenberg limit (HL) Giovannetti et al. (2011); Tóth and Apellaniz (2014); Nawrocki (2019); Polino et al. (2020), where the uncertainty decreases with $1/N$, consists of a well-prepared probe system (in e.g., the NOON state), a definite unitary transformation to encode the phase parameter $\theta$ into the probe state, a practical measurement, and a suitable data processing to produce an estimator $\tilde{\theta}(x)$ about the parameter $\theta$ from the outcome $x$. The estimation precision is quantified by the standard deviation $\delta\theta=\langle[\tilde{\theta}(x)-\theta]^{2}\rangle^{1/2}$, which is under the constraint of the Cramér-Rao inequality Helstrom (1969); Holevo (1982): $\delta\theta\geq 1/\sqrt{F_{c}}$ with $F_{c}$ the classical Fisher information (CFI). Maximizing $F_{c}$ over all possible measurements gives rise to the quantum Fisher information (QFI) $F_{Q}$ and hence the quantum Cramér-Rao bound $\delta\theta_{\rm min}=1/\sqrt{F_{Q}}$ Braunstein and Caves (1994); Braunstein et al. (1996); Luo (2003); Pezzé and Smerzi (2009); Kacprowicz et al. (2010) on the attainable sensitivity to estimate $\theta$.
The optimal precision $\delta\theta_{\rm min}$ generally improves with an increasing quantum number $N$ of the probe state employed in the measurement. With the probe system in separable states, we have $F_{Q}\sim N$ for the standard quantum limit. With particular entangled probe states, such as the NOON state, we have the so-called Heisenberg scaling $F_{Q}\sim N^{2}$ and hence the metrological precision attains HL Leibfried et al. (2004); Mitchell et al. (2004); Boto et al. (2000); Gerry (2000). However, the difficulty in obtaining NOON states with a large $N$ hinders its practical applications Bollinger et al. (1996); Walther et al. (2004); Afek et al. (2010); Zhang et al. (2013). An alternative method to improve the precision is to use the squeezed spin state Wineland et al. (1992); Kitagawa and Ueda (1993); Wineland et al. (1994); Goda et al. (2008), that is typically generated by the one-axis twisting (OAT) and two-axis twisting (TAT) interactions Liu et al. (2011); Zhang et al. (2017). The collective OAT interaction $H_{\rm OAT}\propto J_{z}^{2}$ gives rise to a squeezing degree $\propto 1/N^{2/3}$ for $N$ particles, which has been implemented in Bose-Einstein condensation (BEC) systems Gross et al. (2010); Riedel et al. (2010), trapped ions Bohnet et al. (2016); Lu et al. (2019), superconducting qubits Song et al. (2019); Xu et al. (2020), and warm vapors in glass cells Hosten et al. (2016); Bao et al. (2020); Duan et al. (2024). Under the TAT interaction $H_{\rm TAT}\propto J_{x}^{2}+J_{y}^{2}$, that is mainly reported in theoretical schemes Zhang et al. (2017); Borregaard et al. (2017), the squeezing degree $\propto 1/N$ approaches HL. It remains however a challenge in practice since the interaction Hamiltonian is rarely found in current platforms Helmerson and You (2001); Borregaard et al. (2017); Macrì et al. (2020). In quantum optics, the archetype of a metrology experiment for phase estimation is the Mach-Zehnder interferometer. When the beam splitters are replaced with optical parametric amplifiers, the resulting $SU(1,1)$ interferometer can reach HL only by inputting coherent states Yurke et al. (1986); Li et al. (2014, 2016); Jing et al. (2011); Hudelist et al. (2014); Du et al. (2022). However, its sensitivity to the internal losses eventually makes it work even below SQL Marino et al. (2012).
Recently, the coupling with an ancillary system is found to outperform SQL in the estimation precision about the phase parameter. In a protocol for measuring the light rotation with indefinite time direction Xia et al. (2024), an ancillary system serves as a quantum switch leading to a superposition of opposite rotations of the probe system. The rotation precision can thus be enhanced to HL even by using classical probe states. And a precision of $12.9$ nrad on light rotation measurement has been experimentally realized with a $150$-order Laguerre-Gaussian beam. Motivated by the time reversal strategy Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024), based on which a quantum metrology protocol can approach the Cramér-Rao bound for certain parameters, Ref. Luo and Yu (2023) proposed a controlled protocol based on a two-photonic-mode system and an ancillary qubit. In this protocol, the interaction between the photonic system and the qubit can be manipulated by the external control fields to implement the unitary transformations along the forward- and reverse-time directions. The forward-time evolution generates a photonic entangled NOON state to improve QFI and the reverse-time evolution allows CFI to saturate with its quantum counterpart.
In this work, we present a qubit-assisted metrology protocol to estimate the phase parameter $\theta$ encoded into the probe system (a spin ensemble) during a time evolution, which is based on the time-reversal strategy and a projective measurement. With a properly prepared ancillary qubit, QFI can be modified from the variance of the phase generator to the mean square of that with respect to the probe state. In this case, QFI can be enhanced up to the Heisenberg scaling in terms of the total spin number ($F_{Q}\propto N^{2}$) even when the probe spin ensemble is prepared as a classical state. The scaling behavior is robust against the imperfections about the initial state of the ancillary qubit and the optimized evolution time of the whole system. Moreover, CFI of our protocol is found to be coincident with QFI by the time-reversal strategy.
The rest of this work is structured as follows. In Sec. II, we describe the circuit model of our metrology protocol with the ancillary qubit. In Sec. III, we investigate the conditions for the time-reversal strategy. A detailed derivation about the relevant evolution operator is provided in Appendix A. In Sec. IV, we find that QFI in our protocol can be expressed as the mean square of an optimized phase generator with respect to the probe initial state. That underlies the Heisenberg-limited sensitivity for a large $N$ even when the probe is in the mixed state. In Sec. V, we calculate CFI in our protocol as the extractable information from the probability distribution of the output state upon the projective measurement. In Sec. VI, we adapt our protocol to an alternative interaction between the probe system and an ancillary qubit. We summarize our work in Sec. VII.
II Hamiltonian and metrology model
Consider a model consisting of a spin ensemble as a measurement probe and an ancillary spin-$1/2$. The full Hamiltonian including the free Hamiltonian $H_{0}$ and interaction part $H_{I}$ reads ($\hbar\equiv 1$)
$H=H_{0}+H_{I}=\omega_{P}J_{z}+\omega_{A}\sigma_{z}+gJ_{z}\sigma_{z},$ (1)
where $J_{\alpha}=\sum_{k=1}^{N}\sigma_{\alpha}^{k}/2$, $\alpha=x,y,z$, denotes the collective spin operator with $\sigma_{\alpha}^{k}$ the Pauli matrix of the $k$th probe spin, $\omega_{P}$ and $\omega_{A}$ represent the energy splitting of the probe spin and the ancillary spin, respectively, and $g$ is the coupling strength between the two components.
The full Hamiltonian (1) is feasible in various experiments. In quantum dots Gillard et al. (2022), $J_{z}$ and $\sigma_{z}$ describe the nuclear spins and the electron spin, respectively. In BEC, the Hamiltonian can be realized by two cavities containing two-component BECs coupled by an optical fiber Pyrkov and Byrnes (2013), where the operators $J_{z}$ and $\sigma_{z}$ correspond to the Schwinger boson operators $a^{\dagger}a-b^{\dagger}b$ with $a^{\dagger}$ and $b^{\dagger}$ being bosonic creation operators for two orthogonal quantum states. Also, our interaction Hamiltonian is similar to that by coupling an atomic ensemble to a light beam with the off-resonant Faraday interaction Bao et al. (2020); Duan et al. (2024), where $J_{z}$ is the sum of the total angular momentum of individual atoms and $\sigma_{z}$ denotes the Stokes operator associated with the distinction between the number operators of the photons polarized along orthogonal bases.
Figure 1: Circuit model of our qubit-assisted metrology consisting of a two-step evolution interrupted by the parametric encoding and a single-shot measurement. Step $1$ and $2$ denote the free joint unitary evolution $U(t_{1})$ and its time reversal $U(t_{2})$, respectively, in between which a to-be-estimated parameter $\theta$ is encoded into the probe state via a unitary operation $R_{x}(\theta)$. The output state is determined by a projective measurement on either the full system or the ancillary qubit.
The probe spin ensemble and the ancillary qubit are initially separable, i.e., the input state of the full system is a product state $\rho_{P}\otimes\rho_{A}$. The whole circuit in Fig. 1 can be described by
$U_{\theta}=U(t_{2})R_{x}(\theta)U(t_{1})=e^{-iHt_{2}}e^{-i\theta J_{x}}e^{-iHt_{1}},$ (2)
where the two steps of evolution last $t_{1}$ and $t_{2}$, respectively. During Step $1$, the probe and ancillary system are entangled. Through the parametric encoding of a negligible evolution time, we obtain a phase parameter $\theta$ by a spin rotation $R_{x}(\theta)=\exp(-i\theta J_{x})$ of the probe system. Experimentally $\theta$ can be accumulated during the precessing about $z$-axis Meyer et al. (2001); Gross et al. (2010); Ockeloen et al. (2013) induced by a certain interaction between the probe and a to-be-measured system. Then the rotation about $x$-axis could be performed by a sequence of $R_{x}(\theta)=R_{y}(\pi/2)R_{z}(\theta)R_{y}(-\pi/2)$, where $R_{y}(\pm\pi/2)$ indicates a $\pi/2$ pulse about $y$-axis. The duration time of Step $2$ is determined by $t_{1}$ under the time-reversal strategy Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024). On the last stage of the circuit, a projective measurement about the full system is performed on the output state $\rho(\theta)\equiv U_{\theta}\rho_{P}\otimes\rho_{A}U_{\theta}^{\dagger}$, whose succuss probability can be used to infer the Fisher information about the estimated parameter. When the input state of the probe (spin ensemble) is an eigenstate of $U_{\theta}$, the maximum parameter information can be obtained only by performing a projective measurement on the ancillary qubit.
It has been shown Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024) that the measurement precision can be certainly enhanced such that CFI is increased to be equivalent to QFI when the joint evolution operator $U(t_{2})$ becomes the time reversal of $U(t_{1})$, i.e.,
$U(t_{2})=U^{\dagger}(t_{1}).$ (3)
It indicated that the full system traces back its entanglement generation trajectory and returns to the input state if no phase is encoded.
III Conditions for time reversal
According to the definition of $U_{\theta}$ in Eq. (2) and the requirement for time reversal in Eq. (3), it is evident that
$U_{\theta=0}=e^{-iH(t_{1}+t_{2})}=U(t_{1}+t_{2})=\mathcal{I}^{2(N+1)}$ (4)
$=\sum_{m=-j}^{j}|j,m\rangle\langle j,m|\otimes(|e\rangle\langle e|+|g\rangle\langle g|),$
where $N=2j$ is the total spin number of the probe ensemble, $\mathcal{I}^{n}$ is the identity matrix with dimension $n$, $|j,m\rangle$ is the eigenstate of the operators $J_{z}$ with eigenvalue $m$, and $|g\rangle$ and $|e\rangle$ denote the ground and excited states of the ancillary spin, respectively. Using the full Hamiltonian (1), we have
$U(t_{1}+t_{2})$ $=U(T)=\sum_{m=-j}^{j}|j,m\rangle\langle j,m|$ (5)
$\otimes(C_{e,m}|e\rangle\langle e|+C_{g,m}|g\rangle\langle g|),$
where $C_{e,m}=\exp\{-i[\omega_{A}+m(\omega_{P}+g)]T\}$ and $C_{g,m}=\exp\{i[\omega_{A}-m(\omega_{P}-g)]T\}$ with $T\equiv t_{1}+t_{2}$. It is found that when
$C_{e,m}=C_{g,m^{\prime}},$ (6)
Eq. (4) becomes equivalent to Eq. (5) up to a global phase. Given the expression in Eq. (6), we have
$\left(e^{igT}\right)^{m+m^{\prime}}\left(e^{i\omega_{P}T}\right)^{m-m^{\prime}}=e^{-2i\omega_{A}T}.$ (7)
The parity of the probe-spin number $N$ determines whether the eigenvalue $m$ is an integer or a half-integer. When $N$ is even, $m$ is an integer and then $m\pm m^{\prime}$ are the same in parity. A sufficient solution for Eq. (7) is thus $e^{igT}=e^{i\omega_{P}T}=\pm 1$. Straightforwardly we have
$\frac{gT}{\pi}=n_{1},\quad\frac{\omega_{P}T}{\pi}=n_{1}+2n_{2},\quad\frac{\omega_{A}T}{\pi}=n_{3},$ (8)
where $n_{j}$’s, $j=1,2,3$, are proper integers subject to the given magnitudes of $g$, $\omega_{P}$, and $\omega_{A}$. When $N$ is odd, $m\pm m^{\prime}$ are different in parity. The solution can thus be written as
$\frac{gT}{\pi}=n_{1},\quad\frac{\omega_{P}T}{\pi}=n_{1}+2n_{2},\quad\frac{\omega_{A}T}{\pi}=\frac{n_{1}}{2}+n_{4},$ (9)
where $n_{j}$’s, $j=1,2,4$, are proper integers. For either even or odd, it might be hard to find the exact solution of $T$ if $g$, $\omega_{P}$, and $\omega_{A}$ are incommensurable.
Figure 2: Normalized trace $\mathcal{F}(T)$ as a function of the full evolution time $T$ with various spin number $N$. $\omega_{P}=\omega_{A}=3g$.
The analytical results about $T$ can be confirmed by the numerical simulation over the normalized trace of the unitary operator $U(T)$,
$\mathcal{F}(T)=\frac{|{\rm Tr}[U(T)]|}{2(N+1)}.$ (10)
When $\mathcal{F}(T)=1$, we have $U(T)=\mathcal{I}^{2(N+1)}$ up to a global phase and vise versa. Once $T$ is specified by a unit trace, the time-reversal condition in Eq. (3) is automatically satisfied by setting $t_{2}=T-t_{1}$.
In Fig. 2, we demonstrate the normalized trace $\mathcal{F}(T)$ as a function of the evolution time $T$ exemplified with $N=4$ and $N=11$ under a fixed setting that $\omega_{P}=\omega_{A}=3g$. The parity of the probe-spin number $N$ determines the time when the normalized trace $\mathcal{F}(T)$ attains unit. When $N=4$ (the even case, see the blue solid line), it is found that the normalized trace attains the maximum value when the full evolution time $T$ is a multiple of $\pi/g$. For example, when $T=\pi/g$, Eq. (8) is valid for $n_{1}=1$, $n_{2}=1$, and $n_{3}=3$. When $N=11$ (the odd case, see the red dashed line), the normalized trace attains unit when $T$ is a multiple of $2\pi/g$. Particularly, $T=2\pi/g$ is attained by setting $n_{1}=2$, $n_{2}=2$, and $n_{4}=5$ in Eq. (9). Moreover in case of an odd $N$, one can find that the unitary operator becomes
$U(T)=\mathcal{I}^{N+1}\otimes\sigma_{z}$ (11)
up to a global phase when $T$ is an odd multiple of $\pi/g$. It will give rise to a vanishing normalized trace $\mathcal{F}(T)=0$. And then the time-reversal strategy is broken. Therefore, choosing a probe ensemble with an even number of spins can significantly accelerate our metrology protocol.
IV Quantum Fisher information
Suppose that we have found an exact solution about $T$ on demand of the time-reversal strategy, then the whole unitary evolution operator in Eq. (2) can be rewritten as
$U_{\theta}=e^{iHt_{1}}e^{-i\theta J_{x}}e^{-iHt_{1}}.$ (12)
Using the Baker-Campbell-Hausdorff formula, we have (the details are provided in Appendix A)
$U_{\theta}=e^{-i\theta\left[\cos(gt_{1})J(-\phi)-\sin(gt_{1})J\left(\frac{\pi}{2}-\phi\right)\sigma_{z}\right]},$ (13)
where the phase generator forms as $J(\phi)\equiv\cos(\phi)J_{x}+\sin(\phi)J_{y}$ with $\phi=\omega_{P}t_{1}$. We are now on the stage of analyzing the precision limit of the qubit-assisted metrology.
Initially, the ancillary qubit is assumed to be a pure qubit $\rho_{A}=|\varphi\rangle\langle\varphi|$ with Kitagawa and Ueda (1993); Jin et al. (2009)
$|\varphi\rangle=\cos\left(\frac{\theta_{0}}{2}\right)|e\rangle+e^{-i\phi_{0}}\sin\left(\frac{\theta_{0}}{2}\right)|g\rangle.$ (14)
Here $\theta_{0}$ and $\phi_{0}$ determine the population imbalance and the relative phase between the two bases, respectively. Without loss of generality, we set the azimuthal angle $\phi_{0}=0$. For the input state of the probe system, always it can be written as
$\rho_{P}=\sum_{i=1}^{d}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ (15)
in spectral decomposition, where $d$ is the dimension of the density matrix, $p_{i}$’s are the eigenvalues, and $|\psi_{i}\rangle$’s are the corresponding eigenstates.
Using Eqs. (14) and (15), QFI of the output state associated with the metrology precision for estimating $\theta$ is given by Braunstein and Caves (1994); Braunstein et al. (1996); Zhang et al. (2013); Liu et al. (2014)
$F_{Q}$ $=\sum_{i=1}^{d}4p_{i}\langle\psi_{i}|\langle\varphi|(\partial_{\theta}U_{\theta})^{\dagger}(\partial_{\theta}U_{\theta})|\varphi\rangle|\psi_{i}\rangle$ (16)
$-\sum_{i,j=1}^{d}\frac{8p_{i}p_{j}}{p_{i}+p_{j}}|\langle\psi_{i}|\langle\varphi|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\varphi\rangle|\psi_{j}\rangle|^{2}$
with $p_{i}\neq 0$. We actually obtain a difference between a positive expression, that is vital to achieve Heisenberg scaling as shown in the following, and a non-positive one. A straightforward method to enhance QFI is thus to minimize the magnitude of the second term in Eq. (16) or more generally and precisely to minimize the expression inside the absolute value. Using Eq. (13), we have
$\langle\psi_{i}|\langle\varphi|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\varphi\rangle|\psi_{i}\rangle=\cos(gt_{1})\langle J(-\phi)\rangle_{\psi_{i}}$ (17)
$-\sin(gt_{1})\langle\sigma_{z}\rangle_{\varphi}\left\langle J\left(\frac{\pi}{2}-\phi\right)\right\rangle_{\psi_{i}},$
where $\langle A\rangle_{\psi}\equiv\langle\psi|A|\psi\rangle$. Under a proper state of the ancillary qubit and an optimized evolution time for Step $1$, i.e.,
$\theta_{0}^{\rm opt}=\frac{\pi}{2},\quad t_{1,{\rm opt}}=\left(n+\frac{1}{2}\right)\frac{\pi}{g},$ (18)
with an integer $n$, the two terms in Eq. (17) can vanish at the same time for arbitrary probe state $\psi_{i}$. Consequently, Eq. (13) becomes $U_{\theta}=\exp[i\theta J(\pi/2-\phi_{\rm opt})\sigma_{z}]$ with $\phi_{\rm opt}\equiv\omega_{P}t_{1,{\rm opt}}$.
We emphasize that Eq. (17) presents the main advantage of our qubit-assisted metrology over a conventional parameter estimation Giovannetti et al. (2011); Tóth and Apellaniz (2014); Nawrocki (2019); Polino et al. (2020). For a conventional metrology, in which only the probe state is considered, Eq. (17) becomes
$\langle\psi_{i}|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\psi_{i}\rangle=\langle\cos(\omega_{P}t_{1})J_{x}-\sin(\omega_{P}t_{1})J_{y}\rangle_{\psi_{i}}.$ (19)
Hardly one can find a state-independent time $t_{1}$ leading to $\langle\psi_{i}|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\psi_{i}\rangle=0$, since $\cos(\omega_{P}t_{1})$ and $\sin(\omega_{P}t_{1})$ cannot simultaneously become zero.
Figure 3: (a) QFI as a function of $\theta_{0}$ with a fixed $t_{1}=\pi/(2g)$ and various probe-spin number $N$, (b) QFI as a function of the renormalized evolution time $gt_{1}$ with a fixed $\theta_{0}=\pi/2$ and various $N$, (c) Renormalized QFI $F_{Q}/N^{2}$ in space of $\theta_{0}$ and $gt_{1}$ with $N=4$, and (d) QFI as a function of $N$ for $\rho_{P}=\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$, $\theta_{0}=\pi/2$, and $t_{1}=\pi/(2g)$ (red crosses); for $\rho_{P}^{\rm opt}$, $\theta_{0}=\pi/2$, and $t_{1}=3\pi/(4g)$ (orange squares); for $\rho_{P}^{\rm opt}$, $\theta_{0}=\pi/8$, and $t_{1}=\pi/(2g)$ (green triangles); for $\rho_{P}^{\rm th}$ with $\beta=1$, $\theta_{0}=\pi/2$, and $t_{1}=\pi/(2g)$ [blue circles for numerical results and blue solid line for analytical expression in Eq. (23)]. In (a), (b), (c), and (d), the black dashed line and the black dot-dashed line indicate the Heisenberg scaling $N^{2}$ and the shot-noise scaling $N$, respectively. $\omega_{P}=\omega_{A}=3g$
By Eqs. (16) and (18), QFI becomes the mean square of a phase generator with respect to the probe state, i.e.,
$F_{Q}=4\sum_{i=1}^{d}p_{i}\left\langle J\left(\frac{\pi}{2}-\phi_{\rm opt}\right)^{2}\right\rangle_{\psi_{i}},$ (20)
which is a weighted average over the expectation value for the operator $J(\pi/2-\phi_{\rm opt})^{2}$ within each eigenstate $\{|\psi_{i}\rangle\}$ of $\rho_{P}$. Straightforwardly, QFI can attains its peak value $F_{Q}=N^{2}$ when the probe spin ensemble is prepared in a pure state $|j,j\rangle_{\rm opt}$, or $|j,-j\rangle_{\rm opt}$, and or a mixed state over them, such as $\rho_{P}=(|j,j\rangle_{\rm opt}\langle j,j|+|j,-j\rangle_{\rm opt}\langle j,-j|)/2$, where $|j,m\rangle_{\rm opt}$ with $-j\leq m\leq j$ denotes the eigenstate for the optimized collective angular momentum operator $J(\pi/2-\phi_{\rm opt})$. Subjected to the duration time $t_{1,{\rm opt}}$ and the time-reversal strategy, the bases $|j,m\rangle_{\rm opt}$’s are determined by the magnitudes of $g$, $\omega_{P}$, and $\omega_{A}$. Heisenberg scaling can therefore be approached under optimized $\theta_{0}$ and $t_{1}$.
This condition could be confirmed by the numerical simulation in Fig. 3, where the input state of the probe ensemble is set as $\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$ and the frequencies of both probe spin and ancillary qubit are the same as in Fig. 2. Figure 3(a) describes the dependence of the quantum Fisher information $F_{Q}$ on the input parameter of the ancillary qubit $\theta_{0}$ under an optimized evolution time $t_{1,{\rm opt}}$ given by Eq. (18). It is shown that $F_{Q}$ follows the Heisenberg scaling when $\theta_{0}=\pi/2$ for various $N$. In Fig. 3(b), the initial state of the ancillary qubit is fixed with $\theta_{0}=\pi/2$ and $F_{Q}$ finds its peak at $t_{1}=\pi/(2g)$ as expected by Eq. (18). In both Fig. 3(a) and Fig. 3(b), $F_{Q}$ is symmetrical to the optimized points and varies smoothly around them, indicating that our metrology protocol is not sensitive to the imperfections in parametric control.
In Fig. 3(c), we take $N=4$ and demonstrate the renormalized QFI $F_{Q}/N^{2}$ in the space of $\theta_{0}$ and $t_{1}$. The black dot-dashed lines are the contour lines indicating the proportional coefficients. The central region around point-$A$ describes the most optimized condition that is in agreement with Eq. (18). The surrounding regions about point-$B$ and point-$C$ describe less optimized conditions with $t_{1}$ or $\theta_{0}$ largely departing from Eq. (18). The behaviors for these points with various $N$ are plotted with the red crosses, the orange squares, and the green triangles, respectively, in Fig. 3(d). The first one sticks to its upper bound $F_{Q}=N^{2}$, and the last two are apparently below this bound. However, a similar behavior to Heisenberg scaling $F_{Q}\propto N^{2}$ can be restored when $N$ is sufficiently large, although $\theta_{0}$ or $gt_{1}$ is not optimized.
It is interesting to find that QFI could approach Heisenberg scaling even if the probe ensemble is prepared as a thermal state in the bases of $|j,m\rangle_{\rm opt}$, i.e.,
$\rho_{P}^{\rm th}=\frac{1}{Z_{\beta}}e^{-\beta J(\pi/2-\phi_{\rm opt})},$ (21)
where $Z_{\beta}={\rm Tr}\{\exp[-\beta J(\pi/2-\phi_{\rm opt})]\}$ is the partition function and $\beta$ is the inverse temperature. By Eqs. (20) and (21), we have
$F_{Q}$ $=4\frac{\sum_{m=-N/2}^{N/2}m^{2}\exp(-m\beta)}{\sum_{m=-N/2}^{N/2}\exp(-m\beta)}$ (22)
$=\frac{1}{(e^{\beta}-1)^{2}}\left[N^{2}\frac{1-e^{(3+N)\beta}}{1-e^{(1+N)\beta}}\right.$
$+(N+2)^{2}\frac{e^{2\beta}-e^{(1+N)\beta}}{1-e^{(1+N)\beta}}$
$\left.+(N^{2}+2N-2)\frac{2e^{(2+N)\beta}-2e^{\beta}}{1-e^{(1+N)\beta}}\right].$
And then for a large probe-spin number, i.e., $N\gg 1$, it is approximated as
$F_{Q}$ $\approx\frac{e^{2\beta}N^{2}+(N+2)^{2}-2e^{\beta}(N^{2}+2N-2)}{(e^{\beta}-1)^{2}}$ (23)
$=N^{2}-\frac{4}{e^{\beta}-1}N+4\frac{e^{\beta}+1}{(e^{\beta}-1)^{2}}.$
Clearly $F_{Q}=N^{2}$ when $\beta\rightarrow\infty$. It is indicated that the Heisenberg-scaling behavior $N^{2}$ dominates QFI at least in the low-temperature limit. And $F_{Q}$ will become divergent in the high-temperature limit, i.e., $\beta\rightarrow 0$, which cannot be used to estimate the metrology precision. In Fig. 3(d), the analytical expression in Eq. (23) is confirmed to match the numerical result and both of them approach the Heisenberg scaling, provided $N\gg 1$.
V Classical Fisher information
In a practical scenario of parametric estimation, CFI is defined by catching the amount of information encoded in the probability distribution for the output state under the estimation process Braun et al. (2018); Liu et al. (2020); Tan et al. (2021). And it is upper bounded by its quantum counterpart on the metrology precision. With the time-reversal strategy, it was shown Yurke et al. (1986); Agarwal and Davidovich (2022); Wang et al. (2024) that CFI can be saturated with QFI, which indicates that all the information encoded in the probe state from the output probability distribution has been extracted. We now derive CFI of our protocol under the same optimized settings in Eq. (18) with $n=0$ as for QFI. For the input state $\rho_{P}\otimes\rho_{A}$ indicated by Eqs. (14) and (15), the output state after the two-step evoluation reads
$\rho(\theta)$ $=U_{\theta}\rho_{P}\otimes\rho_{A}U_{\theta}^{\dagger}$ (24)
$=\sum_{m,m^{\prime}=-j}^{j}\sum_{i=1}^{d}\frac{p_{i}p_{i,m,m^{\prime}}}{2}|j,m\rangle_{\rm opt}\langle j,m^{\prime}|$
$\otimes\left(e^{i\theta m}|e\rangle+e^{-i\theta m}|g\rangle\right)\left(e^{-i\theta m^{\prime}}\langle e|+e^{i\theta m^{\prime}}\langle g|\right),$
where $p_{i}$’s are the eigenvalues for the eigenstates $|\psi_{i}\rangle$ of $\rho_{P}$ and $p_{i,m,m^{\prime}}\equiv\langle j,m|\psi_{i}\rangle\langle\psi_{i}|j,m^{\prime}\rangle_{\rm opt}$. Subsequently, we perform the projective measurements $\{|\Psi_{m,\pm}\rangle\langle\Psi_{m,\pm}|\}$ on the full system as described in Fig. 1, where $|\Psi_{m,\pm}\rangle=|j,m\rangle_{\rm opt}\otimes|\pm\rangle$ with |±$|\pm\rangle=(|e\rangle\pm|g\rangle)/\sqrt{2}$, the eigenstates of $\sigma_{x}$ for the ancillary qubit. Note $|j,m\rangle_{\rm opt}$’s are eigenstates of $J(\pi/2-\phi_{\rm opt})$ determined by $g$, $\omega_{P}$, and $\omega_{A}$ in advance. The probability about detecting the output state in the basis $|\Psi_{m,\pm}\rangle$ reads
$P(m,\pm|\theta)$ $=\langle\Psi_{m,\pm}|\rho(\theta)|\Psi_{m,\pm}\rangle$ (25)
$=\frac{1}{4}\sum_{i=1}^{d}p_{i}p_{i,m,m}\left|e^{i\theta m}\pm e^{-i\theta m}\right|^{2}$
$=\frac{1\pm\cos(2m\theta)}{2}\sum_{i=1}^{d}p_{i}p_{i,m,m}.$
By the probability distribution $P(m,\pm|\theta)$ Braun et al. (2018); Liu et al. (2020); Tan et al. (2021), CFI can be calculated as
$F_{c}=\sum_{m=-j}^{j}\left\{\frac{[dP(m,+|\theta)/d\theta]^{2}}{P(m,+|\theta)}+\frac{[dP(m,-|\theta)/d\theta]^{2}}{P(m,-|\theta)}\right\}.$ (26)
The first derivative of $P(m,\pm|\theta)$ with respect to the to-be-estimated phase $\theta$ is $dP(m,\pm|\theta)/d\theta=\mp m\sin(2m\theta)\sum_{i=1}^{d}p_{i}p_{i,m,m}$ and hence
$\frac{[dP(m,\pm|\theta)/d\theta]^{2}}{P(m,\pm|\theta)}=\frac{2m^{2}\sin^{2}(2m\theta)}{1\pm\cos(2m\theta)}\sum_{i=1}^{d}p_{i}p_{i,m,m}.$ (27)
Inserting Eq. (27) to Eq. (26), we have
$F_{c}=$ $\sum_{m=-j}^{j}2m^{2}\left(\sum_{i=1}^{d}p_{i}p_{i,m,m}\right)\sin^{2}(2m\theta)$ (28)
×$\times\left[\frac{1}{1+\cos(2m\theta)}+\frac{1}{1-\cos(2m\theta)}\right]$
$=$ $4\sum_{m=-j}^{j}\sum_{i=1}^{d}m^{2}p_{i}p_{i,m,m}$
$=$ $4\sum_{i=1}^{d}p_{i}\langle\psi_{i}|\left(\sum_{m=-j}^{j}m^{2}|j,m\rangle_{\rm opt}\langle j,m|\right)|\psi_{i}\rangle$
$=$ $4\sum_{i=1}^{d}p_{i}\langle\psi_{i}|J\left(\frac{\pi}{2}-\phi_{\rm opt}\right)^{2}|\psi_{i}\rangle=F_{Q}.$
The last line is exactly the same as Eq. (20) under the optimized condition in Eq. (18), independent of $\rho_{P}$. It is thus verified that QFI can attain its upper-bound in our qubit-assisted metrology. In addition, when the probe is prepared as a polarized state $\rho_{P}^{\rm opt}=|j,\pm j\rangle_{\rm opt}\langle j,\pm j|$, Eq. (25) reduces to
$P(m,\pm|\theta)=\delta_{m,j}\frac{1\pm\cos(2m\theta)}{2},$ (29)
which implies that the magnitude of $\theta$ can be inferred from the probability $P(j,+|\theta)$ and $P(j,-|\theta)$. In this case, only performing the projective measurement on the ancillary qubit $\rho_{A}$ yields the Heisenberg-scaling limit $F_{c}=F_{Q}=N^{2}$.
Figure 4: (a) Renormalized CFI $F_{c}/N^{2}$ in space of $gt_{1}$ and $gt_{2}$. (b) Renormalized CFI $F_{c}/N^{2}$ as a function of $gt_{2}$ with a fixed $gt_{1}=\pi/2$ (blue solid line) and its maximum value obtained with a varying $gt_{1}$ (red dashed line). The vertical black dashed line represents the analytical result in Eq. (9). The probe spin ensemble and the ancillary qubit are initialized in the state $\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$ and $|\varphi\rangle=(|e\rangle+|g\rangle)/\sqrt{2}$, respectively. The other parameters are $\phi_{\rm opt}=\omega_{P}\pi/(2g)$, $N=5$, and $\omega_{P}=\omega_{A}=3g$.
[…]
It is interesting to find that CFI also attains the Heisenberg scaling $F_{c}=N^{2}$ when $t_{2}=\pi/(2g)$ [see the vertical black dot-dashed line in Fig. 4(b)]. It is beyond the sufficient condition for the time-reversal strategy as illustrated by Eq. (9) and Fig. 2. However, it is straightforward to see that when $t_{1}=t_{2}=\pi/(2g)$, the full time-evolution operator in the end can be written as $U^{\prime}_{\theta}=U(\pi/g)U^{\dagger}(t_{1}){R}_{x}(\theta)U(t_{1})=\sigma_{z}U_{\theta}$, where $U_{\theta}$ is given by Eq. (13). Then using Eq. (24), we have $\rho^{\prime}(\theta)=\sigma_{z}U_{\theta}\rho_{P}\otimes\rho_{A}U_{\theta}^{\dagger}\sigma_{z}=\sigma_{z}\rho(\theta)\sigma_{z}$ and hence the probability distribution in Eq. (29) becomes $P^{\prime}(j,\pm|\theta)=[1\mp\cos(2m\theta)]/2$, which means $P^{\prime}(j,\pm|\theta)=P(j,\mp|\theta)$. It indicates that one can still extract all the information encoded in the probe state from the output probabilities and thus $F_{c}=N^{2}$.
VI Discussion
Quantum metrology is conventionally regarded as a parametric-estimation protocol using entanglement or quadratical squeezing to improve the estimation precision beyond the standard quantum limit with respect to the system size $N$. A benchmark result is that the photonic entangled NOON states are capable to attain the Heisenberg scaling characterized with an $N^{2}$-dependent quantum Fisher information. However, a feasible implementation is limited to the number of photons or excitations due to the difficulty in generating a large-scale and high-fidelity NOON state Bollinger et al. (1996); Walther et al. (2004); Afek et al. (2010); Zhang et al. (2013), which achieves up to $N=5$ so far in experiment Afek et al. (2010). As for the spin squeezing that is generated in the nonlinear atom-atom interaction platform, the OAT squeezing demonstrates a scaling below the Heisenberg-limited measurement precision with $\propto 1/N^{2/3}$, e.g., in the cold thermal ensembles Hosten et al. (2016), the metrological gain is then found to be up to $(\delta\theta_{\rm SQL}/\delta\theta)^{2}\sim 100$ with $10^{6}$ atoms.
In contrast, our qubit-assisted metrology protocol can be enhance QFI up to the Heisenberg scaling of $N^{2}$ with respect to the total probe-spin number $N$, even when the probe system is in a mixed state. More than the $ZZ$ interaction between the ancillary qubit and the probe large-spin in the full Hamiltonian (1), our protocol is applicable to another spin-spin-bath model Caldeira et al. (1993); Shao et al. (1996); Xu et al. (2011); Wang and Shao (2012) with Hamiltonian
$H=H_{0}+H_{I}=\omega_{P}J_{z}+\omega_{A}\sigma_{z}+gJ_{x}\sigma_{z}.$ (30)
The corresponding circuit model for the qubit-assisted metrology in Fig. 1 is almost invariant, except that the encoded phase parameter $\theta$ is now obtained by another rotation $R_{z}(\theta)=\exp(-i\theta J_{z})$ of the probe system. Thus the circuit is described by
$U_{\theta}=U(t_{2})R_{z}(\theta)U(t_{1})=e^{-iHt_{2}}e^{-i\theta J_{z}}e^{-iHt_{1}}.$ (31)
Using the modified full Hamiltonian (30), Eq. (5) becomes
$U(t_{1}+t_{2})$ $=U(T)=\sum_{m=-j}^{j}\tilde{C}_{e,m}|j,m\rangle_{\phi}\langle j,m|\otimes|e\rangle\langle e|$ (32)
$+\sum_{m^{\prime}=-j}^{j}\tilde{C}_{g,m^{\prime}}|j,m^{\prime}\rangle_{\phi^{\prime}}\langle j,m^{\prime}|\otimes|g\rangle\langle g|,$
where $\tilde{C}_{e/g,m}=\exp[-i(m\tilde{\omega}\pm\omega_{A})T]$ with $\tilde{\omega}\equiv\sqrt{\omega^{2}_{P}+g^{2}}$. Here $|j,m\rangle_{\phi}$ and $|j,m^{\prime}\rangle_{\phi^{\prime}}$ with eigenvalues $m$ and $m^{\prime}$ denote the eigenstates of the operators $(\omega_{P}J_{z}+gJ_{x})/\tilde{\omega}$ and $(\omega_{P}J_{z}-gJ_{x})/\tilde{\omega}$, respectively. According to $U_{\theta}$ in Eq. (31) and the time-reversal condition in Eq. (3), we have
$U_{\theta=0}$ $=e^{-iH(t_{1}+t_{2})}=U(t_{1}+t_{2})=\mathcal{I}^{2(N+1)}$ (33)
$=\sum_{m=-j}^{j}|j,m\rangle_{\phi}\langle j,m|\otimes|e\rangle\langle e|$
$+\sum_{m^{\prime}=-j}^{j}|j,m^{\prime}\rangle_{\phi^{\prime}}\langle j,m^{\prime}|\otimes|g\rangle\langle g|.$
Equation (32) becomes equivalent to Eq. (33) up to a global phase when
$\tilde{C}_{e,m}=\tilde{C}_{g,m^{\prime}}.$ (34)
A sufficient solution for Eq. (34) is found to be independent of the parity of the probe-spin number $N$, and reads,
$\frac{\tilde{\omega}T}{\pi}=2n_{5},\quad\frac{\omega_{A}T}{\pi}=n_{6},$ (35)
where $n_{j}$’s, $j=5,6$, are proper integers.
With a proper $T$ on the time-reversal strategy, the whole unitary evolution operator in Eq. (31) can be rewritten as
$U_{\theta}=e^{iHt_{1}}e^{-i\theta J_{z}}e^{-iHt_{1}}.$ (36)
Using the Baker-Campbell-Hausdorff formula, we have
$U_{\theta}=e^{i\theta(c_{z}J_{z}+c_{x}J_{x}\sigma_{z}+c_{y}J_{y}\sigma_{z})},$ (37)
where the coefficients of the phase generator are
$c_{z}$ $=-\frac{g^{2}}{\tilde{\omega}^{2}}\cos(\tilde{\omega}t_{1})-\frac{\omega_{P}^{2}}{\tilde{\omega}^{2}},$ (38)
$c_{x}$ $=\frac{g\omega_{P}}{\tilde{\omega}^{2}}[\cos(\tilde{\omega}t_{1})-1],$
$c_{y}$ $=-\frac{g}{\tilde{\omega}}\sin(\tilde{\omega}t_{1}).$
We are now on the stage of minimizing the magnitude of the second term in Eq. (16) to enhance QFI. With Eqs. (37), (14), and (15), we have
$\langle\psi_{i}|\langle\varphi|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\varphi\rangle|\psi_{i}\rangle$ (39)
$=$ $c_{z}\langle J_{z}\rangle_{\psi_{i}}+\langle\sigma_{z}\rangle_{\varphi}\langle c_{x}J_{x}+c_{y}J_{y}\rangle_{\psi_{i}}.$
Under a proper state of the ancillary qubit and an optimized evolution time for Step $1$, i.e.,
$\theta_{0}=\frac{\pi}{2},\quad t_{1}=\tilde{\omega}^{-1}\left[\arccos\left(-\frac{\omega_{P}^{2}}{g^{2}}\right)+2n\pi\right],$ (40)
with an integer $n$, both terms in Eq. (39) can vanish at the same time for arbitrary probe state $\psi_{i}$.
Using Eqs. (16) and (40), QFI becomes the mean square of a phase generator, i.e.,
$F_{Q}=4\sum_{i=1}^{d}p_{i}\left\langle(c_{x}J_{x}+c_{y}J_{y})^{2}\right\rangle_{\psi_{i}}.$ (41)
It reaches the maximum value $F_{Q}=N^{2}$ when the probe spin ensemble is a pure state $|j,j\rangle^{\prime}_{\rm opt}$, or $|j,-j\rangle^{\prime}_{\rm opt}$, and or a mixed state over them. Here $|j,m\rangle^{\prime}_{\rm opt}$ with $-j\leq m\leq j$ denotes the eigenstate for the optimized collective angular momentum operator $c_{x}J_{x}+c_{y}J_{y}$.
[…]
Since $\arccos(x)$ ranges from $-1$ to $1$, the optimal evolution time $t_{1}$ in Eq. (40) can only be obtained in the deep strong coupling regime $g\geq\omega_{P}$. When $g<\omega_{P}$, only the second term in Eq. (39) can vanish for arbitrary probe state and the minimum magnitude of the first term associated with $c_{z}$ can be found at $\cos(\tilde{\omega}t_{1})=-1$. The optimal condition, that does not render $F_{Q}=N^{2}$, can thus be rewritten as
$\theta_{0}=\frac{\pi}{2},\quad t_{1}=\frac{(2n+1)\pi}{\tilde{\omega}},$ (42)
with an integer $n$.
Figure 5: QFI as a function of $N$ for $t_{1}=\pi/(\sqrt{2}g)$ and $g/\omega_{P}=1$ (red crosses); for $t_{1}=\pi/(3.48g)$ and $g/\omega_{P}=0.3$ (blue circles); for $t_{1}=\pi/(\sqrt{101}g)$ and $g/\omega_{P}=0.1$ (orange squares). The orange solid line is a fit curve for the last case with $F_{Q}=aN^{2}+bN$, where $a\approx 0.04$ and $b\approx 0.96$. The black dashed line and the black dot-dashed lines indicate the Heisenberg scaling $N^{2}$ and the shot-noise scaling $N$.
In Fig. 5, we show the dependence of QFI on the probe-spin number $N$ under various $t_{1}$ and $g/\omega_{P}$ according to Eq. (40) or Eq. (42) with $n=0$, where the initial state of the probe ensemble is set as $\rho_{P}^{\prime\rm opt}=|j,j\rangle^{\prime}_{\rm opt}\langle j,j|$. Numerical simulation confirms that when $g=\omega_{P}$, QFI attains its peak value $F_{Q}=N^{2}$. The scaling behavior of QFI gradually deviates from the Heisenberg limit with a decreasing $g$. Yet the numerical fitting for $g/\omega_{P}=0.1$ (see the orange solid line and the orange squares) indicate that QFI will follow the Heisenberg scaling in an asymptotic way. It is found that $F_{Q}\approx 0.04N^{2}+0.96N$, where the two coefficients are dependent on $g/\omega_{P}$ yet independent of $N$.
VII Conclusion
In summary, we incorporate the joint-evolution-and-quantum-measurement idea to quantum parameter estimation in our metrology protocol. With no input of entangled or squeezed state and with no use of nonlinear Hamiltonian, QFI about the phase encoded into the probe system can be optimized to be the mean square of an optimized phase generator with respect to the probe initial state. That renders an exact or asymptotic Heisenberg-scaling behavior in terms of the spin number $N$ even when the probe starts from a thermal state. And the numerical simulation shows that the imperfections in preparing the ancillary qubit and optimizing the evolution time does not change the Heisenberg-scaling behavior about metrology precision as long as $N$ is sufficiently large. By virtue of the time-reversal strategy and the projective measurement on the whole system or merely the ancillary qubit, CFI is found to be saturated with its quantum counterpart. Our protocol of qubit-assisted quantum metrology is applicable to a general spin-spin-bath model. It therefore provides an economical way toward the Heisenberg-scaling metrology.
Acknowledgments
We acknowledge financial support from the National Natural Science Foundation of China (Grant No. 11974311).
Appendix A Evolution operator for time-reversal strategy in the qubit-assisted metrology
This appendix derives the unitary evolution operator for the time-reversal strategy in our qubit-assisted metrology. With the time-reversal condition (3), the whole evolution operator in Eq. (12) can be written as
$U_{\theta}$ $=e^{iHt_{1}}e^{-i\theta J_{x}}e^{-iHt_{1}}$ (43)
$=e^{iHt_{1}}\left[\sum_{k=0}^{\infty}\frac{1}{k!}\left(-i\theta J_{x}\right)^{k}\right]e^{-iHt_{1}}$
$=\sum_{k=0}^{\infty}\frac{1}{k!}\left(-i\theta e^{iHt_{1}}J_{x}e^{-iHt_{1}}\right)^{k}$
$=\exp\left(-i\theta e^{iHt_{1}}J_{x}e^{-iHt_{1}}\right).$
Using the full Hamiltonian (1) and the commutation relation $[J_{z},\sigma_{z}]=0$,
$e^{iHt_{1}}J_{x}e^{-iHt_{1}}$ (44)
$=$ $e^{igt_{1}J_{z}\sigma_{z}}\left(e^{i\omega_{P}t_{1}J_{z}}J_{x}e^{-i\omega_{P}t_{1}J_{z}}\right)e^{-igt_{1}J_{z}\sigma_{z}}$
$=$ $\cos(\omega_{P}t_{1})e^{igt_{1}J_{z}\sigma_{z}}J_{x}e^{-igt_{1}J_{z}\sigma_{z}}$
-- $\sin(\omega_{P}t_{1})e^{igt_{1}J_{z}\sigma_{z}}J_{y}e^{-igt_{1}J_{z}\sigma_{z}}.$
In the last step, we have used $e^{i\alpha J_{z}}J_{x}e^{-i\alpha J_{z}}=\cos(\alpha)J_{x}-\sin(\alpha)J_{y}$. The Baker-Campbell-Hausdorff formula yields:
$e^{igt_{1}J_{z}\sigma_{z}}J_{x}e^{-igt_{1}J_{z}\sigma_{z}}=\sum_{k=0}^{\infty}\frac{1}{k!}C_{k},$ (45)
where $C_{k+1}=igt_{1}[J_{z}\sigma_{z},C_{k}]$ and $C_{0}=J_{x}$. Using the commutation relation $[J_{\alpha},J_{\beta}]=i\epsilon_{\alpha\beta\gamma}J_{\gamma}$ with $\epsilon_{\alpha\beta\gamma}$ the Levi-Civita tensor, we have
$C_{1}$ $=igt_{1}[J_{z}\sigma_{z},C_{0}]=-gt_{1}J_{y}\sigma_{z},$ (46)
$C_{2}$ $=igt_{1}[J_{z}\sigma_{z},C_{1}]=-(gt_{1})^{2}J_{x},$
$C_{3}$ $=igt_{1}[J_{z}\sigma_{z},C_{2}]=(gt_{1})^{3}J_{y}\sigma_{z},$
and so on. Equation (45) thus becomes
$e^{igt_{1}J_{z}\sigma_{z}}J_{x}e^{-igt_{1}J_{z}\sigma_{z}}$ (47)
$=$ $J_{x}\sum_{k=0}^{\infty}\frac{(-1)^{k}}{(2k)!}(gt_{1})^{2k}-J_{y}\sigma_{z}\sum_{k=0}^{\infty}\frac{(-1)^{k}}{(2k+1)!}(gt_{1})^{2k+1}$
$=$ $\cos(gt_{1})J_{x}-\sin(gt_{1})J_{y}\sigma_{z}.$
And similarly, we have
$e^{igt_{1}J_{z}\sigma_{z}}J_{y}e^{-igt_{1}J_{z}\sigma_{z}}=\cos(gt_{1})J_{y}+\sin(gt_{1})J_{x}\sigma_{z}.$ (48)
Therefore, Eq. (44) can be reduced to
$e^{iHt_{1}}J_{x}e^{-iHt_{1}}$ (49)
$=$ $\cos(gt_{1})[\cos(\omega_{P}t_{1})J_{x}-\sin(\omega_{P}t_{1})J_{y}]$
-- $\sin(gt_{1})[\cos(\omega_{P}t_{1})J_{y}+\sin(\omega_{P}t_{1})J_{x}]\sigma_{z},$
and hence
$U_{\theta}=e^{-i\theta\left[\cos(gt_{1})J(-\phi)-\sin(gt_{1})J\left(\frac{\pi}{2}-\phi\right)\sigma_{z}\right]},$ (50)
where $J(\phi)\equiv\cos(\phi)J_{x}+\sin(\phi)J_{y}$ and $\phi=\omega_{P}t_{1}$. It is exactly Eq. (13) in the main text.